OpenStudy (anonymous):

solve 5^x^2-5x = 5^4-2x

OpenStudy (anonymous):

I need to find the value of x

OpenStudy (anonymous):

x=-1, and x=4

OpenStudy (anonymous):

solve 5^x^2-5x = 5^4-2x is that like this? 5^[x^2-5x] = 5^[4-2x] then x^2-5x=4-2x x^2-5x+2x-4=0 x^2-3x-4=0 can you do from here

OpenStudy (anonymous):

thank so much :)

OpenStudy (anonymous):

yw